# Model for Peanuts: Hijacking ML models without Training Access is Possible

### v1

@@ Abstract.
The massive deployment of Machine Learning (ML) models has been accompanied by the emergence of several attacks that threaten their trustworthiness and raise ethical and societal concerns such as invasion of privacy, discrimination risks, lack of accountability, and wariness of unlawful surveillance. Model hijacking is one of these attacks, where the adversary aims to hijack a victim model to execute a different task than its original one. Model hijacking can cause accountability and security risks since a hijacked model owner can be framed for having their model offering illegal or unethical services. Prior state-of-the-art works consider model hijacking as a training time attack, whereby an adversary requires access to the ML model training to execute their attack. In this paper, we consider a stronger threat model where the attacker has no access to the training phase of the victim model. Our intuition is that ML models, typically over-parameterized, might be able to (unintentionally) learn more than the intended task for they are trained. We propose a simple approach for model hijacking at inference time named SnatchML  to classify unknown input samples using distance measures in the latent space of the victim model to previously known samples associated with the hijacking task classes. SnatchML  empirically shows that benign pre-trained models can execute tasks that are semantically related to the initial task. Surprisingly, this can be true even for hijacking tasks unrelated to the original task. We also explore different methods to mitigate this risk. We first propose a novel approach we call meta-unlearning, designed to help the model unlearn a potentially malicious task while training on the original task dataset. We also provide insights on over-parametrization as one possible inherent factor that makes model hijacking easier, and we accordingly propose a compression-based countermeasure against this attack. We believe this work offers a previously overlooked perspective on model hijacking attacks, featuring a stronger threat model and higher applicability in real-life contexts.
Machine Learning, Security, Privacy, Model Hijacking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software and application security
1. Introduction
@@
Machine Learning models have demonstrated cutting-edge performance across a broad spectrum of applications, progressively expanding to be deployed into domains with security-critical and privacy-sensitive implications, such as healthcare, financial sectors, transportation systems, and surveillance. However, as the massive adoption of ML models continues to rise, a variety of attacks with different threat models have emerged, which can jeopardize ML models’ trustworthiness. For example, adversarial attacks  (Carlini and Wagner, 2017; Huang et al., 2011; Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Venceslai et al., 2020; Guesmi et al., 2023) are popular inference time attacks that compromise the security of the model, by causing it to misclassify to the attacker’s advantage. The necessity of large amounts of data and high computational resources at the training stage has introduced another attack surface against ML, where the adversary interferes with the model training. Such attacks are usually referred to as training time attacks. Within this category, backdoor and data poisoning attacks are two of the most popular ones (Sun et al., 2019; Bagdasaryan et al., 2020; Biggio et al., 2012; Naseri et al., 2020).
Model hijacking. Recently, a new type of training time attack known as model hijacking has been introduced (Si et al., 2023; Salem et al., 2022; Elsayed et al., 2018; Mallya and Lazebnik, 2018). In model hijacking attacks, the adversary aims to take control of a target model and repurpose it to perform a completely different task, referred to as hijacking task. Salem et al. (Salem et al., 2022) proposed a Model Hijacking attack that hides a model covertly while training a victim model. Elsayed et al. (Elsayed et al., 2018) proposed adversarial reprogramming in which, instead of creating adversarial instances, they crafted inputs that would trick the network into performing new tasks. In the same direction, Mallya et al. (Mallya and Lazebnik, 2018) proposed Packnet that trains the model with multiple tasks. The conventional execution of these attacks is concretely performed through data poisoning aimed to repurpose the victim model trained for an original task to perform a hijacking task without reducing the original task’s utility.
@@
In a model hijacking scenario, the attacker repurposes a model to perform a malicious task without the model owner’s knowledge or consent. This can cause accountability risks for the model owner; they can be framed as offering illegal or unethical services. For example, an adversary can repurpose a benign image classifier into a facial recognition model for illegal surveillance (e.g., on data powered by IoT bots) or for NSFW movies. Other possible unethical or illegal scenarios include hijacking a benign model to enable systematic discrimination based on gender, race, or age.
The threat model of existing hijacking attacks is similar to data poisoning/backdoor attacks, i.e., it requires access to the training process of the victim model. In this work, we consider an even more critical threat model for model hijacking where the attacker has limited access capabilities. Specifically, we consider a trusted model provider who (securely) trains a benign ML model for an original task. At training time, the adversary cannot access the training data or process. We consider the scenario where the attacker has access to the model at inference time only, i.e., after deployment. Under these assumptions, we ask the following question:
Can a model securely trained for an original task $\mathcal{T}$ be repurposed for a different task $\mathcal{T^{\prime}}$ without retraining?
Figure 1. An attacker can use the output of a model trained on task $\mathcal{T}$ to infer a different task ($\mathcal{T}^{\prime}$). For example, a user can be identified by simply comparing the similarity between the feature maps of a query and a reference identity class.
@@
Considering this threat model, we propose SnatchMLto hijack an ML model with inference time-only access. SnatchML  leverages the capacity of benign ML models trained on clean datasets, using a conventional training process, to acquire ”extra-knowledge” to infer an extra (potentially malicious) task. Having access to the deployed model, the adversary utilizes ”benign extracted knowledge” to infer the hijacking task. Specifically, we analyze the use of either logits (in a black-box scenario) or feature vectors (in a white-box scenario) to classify unknown input samples, using distance measures in the latent space, to previously known samples associated with the classes relevant to the hijacking task. The proposed approach is detailed in Section 4. Figure 1 gives a high-level illustration of the attack setting in a scenario where the original task is emotion recognition, and the hijacking task is biometric identification.
To demonstrate our attack methodology, we analyze various scenarios. This analysis initially focuses on hijacking tasks semantically related to the original task. Specifically, we examine three original tasks where we consider a pre-trained model and demonstrate that, in each case, an attacker with restricted access can exploit the pre-trained model for a hijacking task that shares semantic overlap with the original task (Section 5, 6 and 7).
@@
While these attacks illustrate the hijacking risk under a stronger threat model than what is currently established in the state-of-the-art, the prerequisite of relatedness between the original and hijacking tasks may limit the scope and impact of such attack strategy. Therefore, in Section 8, we investigate the general case where the relatedness constraint is relaxed. Surprisingly, we found that SnatchML  is capable of hijacking a deployed model for a task that is totally unrelated to the original one. We attempt to provide an explanation of these findings in Section 9. We hypothesize that the overparametrization of ML models is a core reason behind this phenomenon; (i) it provides the capacity to learn clues useful for related tasks, and (ii) it results in a hyper-dimensional representation of data in the latent space, which acts akin to a random projection block, enabling the inference of unrelated tasks.
We contend that our work also has significant implications for risk-based regulatory frameworks, as it challenges some of their foundational assumptions. In fact, the debate on regulating AI-powered systems has gained global momentum, ultimately exemplified by initiatives like the European Commission’s proposal for a risk-assessment framework known as the EU AI-Act (Commission, [n. d.]). This framework seeks to identify and categorize potential security risks and safety implications based on the nature of the task for which the ML model is trained, i.e., the original task. The reliance on the learned task’s criticality as a risk metric is inevitably tied to the following implicit hypothesis: ”If an ML model is trained on a data distribution $\mathcal{D}$, to learning a task $\mathcal{T}$, it is unlikely to (unintentionally) learn another task $\mathcal{T^{\prime}}$”. Our work shows that it is not sufficient for a benign model to be securely trained on a legitimate original task to guarantee that it will be immune to repurposing for unethical or illegitimate tasks. Therefore, additional safeguards and assurances are necessary.
@@
We propose two methods to mitigate SnatchML’s risk. We first propose meta-unlearning, which helps unlearn the potentially malicious task while learning the original task. The second defense method is based on our study of the over-parametrization’s impact on enabling model hijacking. We formulate an optimization problem to find the most compact model that preserves the accuracy of the original task while being less prone to hijacking attacks.
@@
  • We investigate the risk of ML hijacking attacks with a strong threat model where the attacker cannot access the training data/process. Specifically, we propose SnatchML  to exploit the model’s unintentionally learned/extracted capabilities.
@@
  • We illustrate our study with practical scenarios and show that models trained for benign tasks can be hijacked for unethical use, such as biometric identification gender or ethnicity recognition. Surprisingly, we also find it possible to infer hijacking tasks totally unrelated to the original task.
  • We investigate over-parametrization as one of the potential fundamental reasons behind the capacity of ML models to break the least privilege principle during their training.
@@
  • We propose two approaches that can be deployed to limit the risk of SnatchML; We propose meta-unlearning, a novel meta-learning-based approach that helps the model unlearn the potentially malicious task when training on the original task. We finally use model compression to strictly limit the model’s capacity to the original task.
This work illustrates new risks to ML models. We hope it will draw attention to the principle of least privilege as a core cybersecurity principle that needs to be considered for the training and deploying of trustworthy ML models.
Figure 2. Correlation distribution between output feature maps from pretraine Emotion Recognition and Face Recognition.
2. Intuition and Preliminary Analysis
We posit that the models might learn more than they should in the training process, including learning other tasks. Our intuition is that ML models, perhaps due to over-parametrization, have a by-design capacity that exceeds the minimum necessary capacity to learn the task for which they are trained. While it has been shown in the literature that models can overfit and memorize data, making them vulnerable to certain attacks such as Membership (Ye et al., 2022) and property inference (Ganju et al., 2018) attacks, the risk that ML models (unintentionally) generalize to other tasks is yet to be explored/exploited. In this section, we provide a preliminary analysis to gain insights into the plausibility of this hypothesis. Given a model $f^{(i)}(.)$ trained on a data distribution $(X,Y)\sim\mathcal{D}$, using a loss function $\mathcal{L}_{i}$, relevant to a Task $\mathcal{T}_{i}$. Let $f^{(j)}(.)$ be a model with the same architecture as $f^{(i)}(.)$, but is trained on a data distribution $(X’,Y’)\sim\mathcal{D}’$ ’, using a loss function $\mathcal{L}_{j}$, relevant to a Task $\mathcal{T}_{j}\neq\mathcal{T}_{i}$. We ask whether features learnt by $f^{(i)}(.)$ are correlated with those learnt by $f^{(j)}(.)$? We train two ResNet-9 and two Mobilenet architectures each for Emotion Recognition and Face Recognition. We then check the correlation coefficient distribution between the feature maps of the same layers that correspond to the 2 tasks. The correlation coefficient between two variables $X$ and $Y$, denoted as $r$, is defined as:
(1) $r=\frac{\sum_{i=1}^{n}(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum_{i=1}^{n}(X_{i}-\overline{X})^{2}}\sqrt{\sum_{i=1}^{n}(Y_{i}-\overline{Y})^{2}}}$
where:
  • $n$ is the number of data points,
  • $X_{i}$ and $Y_{i}$ are the individual values of the variables $X$ and $Y$,
  • $\overline{X}$ and $\overline{Y}$ are the means of $X$ and $Y$ respectively.
Figure 2 shows the features correlation distribution for Layers 5 and 11 for ResNet-9 and Layers 2 and 4 for Mobilenet. Interestingly, Figure 2 shows a positive correlation, suggesting a potential overlap between what the model learns for tasks that share semantic clues. In other words, this preliminary observation supports the idea that models trained on an original task can extract features relevant to other tasks.
3. Threat Model
@@
We consider an attacker who deploys a pre-trained ML model acquired from a trusted third party (vendor). The model is securely trained on clean data to execute a benign original task.
@@
Attacker’s objective: This threat model focuses on scenarios where the attacker aims to re-purpose the deployed ML model for eventual malicious tasks different from the original one. For example, an attacker wants to hijack a compliant pretrained model for unethical or illegal activities without having access to the training.
@@
Attacker’s Capabilities: (i) At training time, We assume that the model is trained to perform a benign task and that the attacker cannot interfere by any means with the model’s training phase, i.e., the adversary cannot poison the target model’s training dataset, in contradiction with existing hijacking (Salem et al., 2022; Si et al., 2023) or poisoning (Biggio et al., 2012) attacks’ assumptions.
@@
(ii) At inference time, we consider that the attacker can have access to the model under two different settings:
@@
(B) White-box: In this case, the attacker can have access to the internal state of the model (e.g., feature maps). This case corresponds to where the model owner is the attacker; they can show the ML model’s compliance with the regulations by referring to the trustworthy vendor but use the model for potentially unethical tasks.
@@
4. SnatchML: General Approach
Problem Statement– Given a model $h_{\theta}(\cdot)$ with parameters $\theta$, securely trained on a clean data distribution $D(X,Y)$ using a Loss function $\mathcal{L}$ to perform a task $\mathcal{T}$. The objective is to leverage the hidden capabilities that the victim model $h_{\theta}(\cdot)$ acquired through the design and training process to exploit the model to the adversary’s advantage post-deployment. Specifically, we want to investigate the following questions:
  • Q1– Can $h_{\theta}(\cdot)$ unintentionally learn information that allows an attacker to hijack it for another task $\mathcal{T^{\prime}}\neq\mathcal{T}$?
  • Q2– Ultimately, can $h_{\theta}(\cdot)$ still be used to infer a task $\mathcal{T^{\prime}}$ that is totally unrelated to $\mathcal{T}$?
Proposed approach– Without loss of generality, let us suppose that the model $h_{\theta}(\cdot)$ is trained to perform a given multi-class classification (original) task $\mathcal{T}$ with $n$ classes. Given the threat model in Section 3, the attacker wants to use the model for another (hijacking) task $\mathcal{T^{\prime}}$.
Definition 4.1.
(Benign Extracted Knowledge). We define the benign extracted knowledge facts (BEK) as metadata learned by a benign model from clean input. For a given model $h(\cdot)$ and an input sample $x$, we note $\zeta_{h}(\cdot)$ an operator that extracts BEK.
Given Definition 4.1, $\zeta_{h}(x)$ may correspond in the black-box scenario to $\zeta_{(}x)=Z_{h}(x)$, where $Z_{h}=\{z_{i}\}_{i=1..n}$ is model’s output logits vector, or in the white-box scenario to the learned features tensor, i.e., $\zeta_{h}(x)=h_{k}(x)$, where $h_{k}(x)$ is the output of the $k^{th}$ layer of the model $h(\cdot)$.
The attacker wants to repurpose $h(\cdot)$ for a hijacking $m$-class classification task $\mathcal{T}^{\prime}$ through post-processing $\zeta_{h}(\cdot)$. The attacker has access to a dataset $\mathcal{D}^{*}={(x^{*}_{i},\ell^{*}_{i}),i\in[1,m]}$, which contains only $m$ data samples, each corresponding to a class of the hijacking task $\mathcal{T}^{\prime}$. We propose a straightforward exploitation method that is based on the distance in the BEK space. Given an input sample $x_{s}$ and the corresponding BEK vector $\zeta_{h}(x_{s})$ after inference with $h(\cdot)$, the classification of $x_{s}$ for $\mathcal{T}^{\prime}$ is inferred as follows:

### v2

# SnatchML: Hijacking ML models without Training Access

@@ Abstract
The widespread deployment of Machine Learning (ML) models has been accompanied by the emergence of various attacks that threaten their trustworthiness and raise ethical and societal concerns. One such attack is model hijacking, where an adversary seeks to repurpose a victim model to perform a different task than originally intended. Model hijacking can cause significant accountability and security risks since the owner of a hijacked model can be framed for having their model offer illegal or unethical services. Prior works consider model hijacking as a training time attack, whereby an adversary requires full access to the ML model training. In this paper, we consider a stronger threat model for an inference-time hijacking attack, where the adversary has no access to the training phase of the victim model. Our intuition is that ML models, which are typically over-parameterized, might have the capacity to (unintentionally) learn more than the intended task they are trained for. We propose SnatchML, a new training-free model hijacking attack, that leverages the extra capacity learnt by the victim model to infer different tasks that can be semantically related or unrelated to the original one. Our results on models deployed on AWS Sagemaker showed that SnatchML can deliver high accuracy on hijacking tasks. Interestingly, while all previous approaches are limited by the number of classes in the benign task, SnatchML  can hijack models for tasks that contain more classes than the original. We explore different methods to mitigate this risk; We propose meta-unlearning, which is designed to help the model unlearn a potentially malicious task while training for the original task. We also provide insights on over-parametrization as a possible inherent factor that facilitates model hijacking, and accordingly, we propose a compression-based countermeasure to counteract this attack. We believe this work offers a previously overlooked perspective on model hijacking attacks, presenting a stronger threat model and higher applicability in real-world contexts. Our code is available at https://github.com/ihsenLab/SnatchML.
I Introduction
@@
Machine Learning models have demonstrated cutting-edge performance across a broad spectrum of applications, progressively expanding to be deployed into domains with security-critical and privacy-sensitive implications, such as healthcare, financial sectors, transportation systems, and surveillance. However, as the massive adoption of ML models continues to rise, a variety of attacks with different threat models have emerged, which can jeopardize ML models’ trustworthiness. For example, adversarial attacks  [9, 23, 18, 44, 20] are popular inference-time attacks that compromise the security of the model by causing it to misclassify to the attacker’s advantage. The necessity of large amounts of data and high computational resources at the training stage has introduced another attack surface against ML, where the adversary interferes with the model training. Such attacks are usually referred to as training time attacks. Within this category, backdoor and data poisoning attacks are two of the most popular ones [43, 3, 7, 33].
@@
Model hijacking. Recently, a new type of training time attacks known as model hijacking has been introduced [40, 38, 13, 30]. In model hijacking attacks, the adversary’s objective is to take control of a target model and repurpose it to perform a completely different task, referred to as hijacking task without the model owner’s knowledge or consent. This can cause accountability risks for the model owner; they can be framed to be offering illegal or unethical services. For example, an adversary can repurpose a benign image classifier into a facial recognition model for illegal surveillance (e.g. on data powered by IoT bots) or for NSFW material. Other possible unethical or illegal scenarios include hijacking a benign model to enable systematic discrimination based on gender, race or age. Among first works on this threat, Salem et al. [38] proposed Model Hijacking attack that hides a model covertly while training a victim model. In the same direction, Mallya et al. [30] proposed Packnet that trains the model with multiple tasks.
The threat model of these attacks is similar to data poisoning/backdoor attacks, i.e., it requires access to the training process of the victim model. However, in this work, we consider an more critical threat model where the attacker has limited access capabilities. Specifically, the adversary has no access to the training data or process and has access to the model at inference time only after deployment. Moreover, while existing hijacking attacks focus on scenarios where the attacker aims to frame the model owner, we also consider a different setting, where the model owner intentionally misuses a benign/compliant model for malicious activities in a covert manner. In this case, they may avoid altering the input data to evade detection. For instance, consider a scenario where a CCTV camera is used by authorities or a company for a legitimate purpose, such as emotion recognition, utilizing a model from a trusted vendor. With SnatchML, the same model can be covertly repurposed on the same input stream for malicious activities, such as tracking minorities.
@@
SnatchML  leverages the capacity of benign ML models trained on clean datasets to acquire ”extra-knowledge” to infer a different (potentially malicious) task. The adversary, having access to the deployed model, utilizes ”benign extracted knowledge” to infer the hijacking task. Specifically, we analyze the use of either logits (in a black-box scenario) or feature vectors (in a white-box scenario) to classify unknown input samples, using distance measures in the latent space, to previously known samples associated with the classes relevant to the hijacking task. The proposed approach is detailed in Section III. Figure 1 gives a high-level illustration of the attack setting in a scenario where the original task is emotion recognition and the hijacking task is biometric identification.
Figure 1: An attacker can use the output of a model trained on task $\mathcal{T}$ to infer a different task ($\mathcal{T}^{\prime}$). For example, a user can be identified by simply comparing the similarity between feature maps of a query and a reference identity class.
@@
To demonstrate our attack methodology, we conduct an analysis across various scenarios with pretrained ML models deployed on AWS. This analysis initially focuses on hijacking tasks that are semantically related to the original task. Specifically, we examine four original tasks where we consider a pre-trained model and demonstrate that, in each case, an attacker with restricted access can exploit a pre-trained model for a hijacking task that shares semantic overlap with the original task (Section IV, V, VI and VII). While these attacks illustrate the hijacking risk under a stronger threat model than what is currently established in literature, the prerequisite of relatedness between the original and hijacking tasks may limit the scope and impact of such attack strategy. Therefore, in Section VIII, we investigate the general case where the relatedness constraint is relaxed. Surprisingly, we found that SnatchML  is capable of hijacking a deployed model for a task that is totally unrelated to the original one. We attempt to provide an explanation of these findings in Section X. Our hypothesis is that the overparametrization of ML models is a core reason behind this phenomenon; (i) it provides capacity to learn clues useful for related tasks, and (ii) it results in a hyper-dimensional representation of data in the latent space, which acts akin to a random projection block, enabling the inference of unrelated tasks.
@@
We propose two methods to mitigate SnatchML’s risk: We first propose meta-unlearning, which helps unlearning the potentially malicious task while learning the original task. The second defense method is based on our study of the over-parametrization’s impact on enabling model hijacking; We formulate an optimisation problem to find the most compact model that preserves the accuracy of the original task while being less prone to hijacking attacks.
@@
  • We propose an inference-time model hijacking attack that does not require training access and has no restrictions on the number of classes for the hijacking task, applicable in both black-box and white-box settings.
@@
  • We illustrate our study with practical scenarios and show that models trained for benign tasks can be hijacked for unethical use such as biometric identification, gender or ethnicity recognition. Surprisingly, we also find that it is possible to infer hijacking tasks that are totally unrelated to the original task.
@@
  • We investigate the over-parametrization as a potential reason behind the vulnerability to inference-time hijacking, and we propose two approaches to limit the risk of SnatchML: (i) Meta-unlearning, a novel meta-learning based approach that helps the model unlearn an identified malicious task when training on the original one, and (ii) Model compression to strictly limit the capacity of the model to the original task.
Disclosures and ethics. We disclosed our findings to Amazon AWS. We tested our experiments on AWS SageMaker [2], specifically using JumpStart, and were limited to our own models. No public or commercial MLaaS systems were impacted.
II Threat Model
@@
We consider an attacker who deploys a pre-trained ML model acquired from a trusted third party. The model is securely trained on clean data for a benign original task.
@@
Attacker’s objective: This threat model focuses on scenarios where the attacker aims to re-purpose the deployed ML model for eventual malicious tasks different from the original one. We consider 2 scenarios: The first corresponds to the classical hijacking where the attacker aims to frame the victim, who is the model owner. The second scenario is when the attacker is the model owner who purposefully misuses a benign/compliant model for malicious activities in a covert manner, i.e., without changing the model or the input.
@@
Attacker’s Capabilities: (i) At training time, We assume that the model is trained to perform a benign task and that the attacker cannot interfere by any means with the model’s training phase, i.e., the adversary does not have the ability to poison the target model’s training dataset, in contradiction with existing hijacking [38, 40] or poisoning [7] attacks’ assumptions.
@@
(ii) At inference time, we consider that the attacker can have access to the model under 2 different settings:
@@
(B) White-box: in this case the attacker can have access to the internal state of the model (e.g., feature maps). This case corresponds to the case where the model owner is the attacker; they can show the ML model’s compliance to the regulations by referring to the trustworthy vendor, but use the model for potentially unethical tasks.
@@
III SnatchML: General Approach
Problem Statement– Given a model $h_{\theta}(\cdot)$ with parameters $\theta$, securely trained on a clean data distribution $D(X,Y)$ using a Loss function $\mathcal{L}$ to perform a task $\mathcal{T}$. The objective is to leverage the hidden capabilities that the victim model $h_{\theta}(\cdot)$ acquired through the design and training process to exploit the model to the adversary’s advantage post-deployment. Specifically, the adversary aims to leverage $h_{\theta}(\cdot)$’s unintentionally learnt information/capabilities to hijack it for another task $\mathcal{T^{\prime}}$.
Proposed approach– Without loss of generality, let’s suppose that the model $h(\cdot)$ is trained to perform an original task $\mathcal{T}$ consisting of a multi-class classification with $n$ classes. Given the threat model in Section II, the attacker wants to use the model for another (hijacking) task $\mathcal{T^{\prime}}$.
Definition 1.
(Benign Extracted Knowledge). We define the benign extracted knowledge facts (BEK) as metadata learnt by a benign model from clean input. For a given model $h(\cdot)$ and an input sample $x$, we note $\zeta_{h}(\cdot)$ an operator that extracts BEK.
Given Definition 1, $\zeta_{h}(x)$ may correspond in the black-box scenario to $\zeta_{(}x)=Z_{h}(x)$, where $Z_{h}=\{z_{i}\}_{i=1..n}$ is model’s output logits vector, or in the white-box scenario to the learnt features tensor, i.e., $\zeta_{h}(x)=h_{k}(x)$, where $h_{k}(x)$ is the output of the $k^{th}$ layer of the model $h(\cdot)$.
The attacker wants to repurpose $h(\cdot)$ for a hijacking $m$-class classification task $\mathcal{T}^{\prime}$ through a 00-shot inference on $\zeta_{h}(\cdot)$. The attacker has access to a dataset $\mathcal{D}^{*}={(x^{*}_{i},\ell^{*}_{i}),i\in[1,m]}$, which contains only $m$ data samples, each corresponding to a class of the hijacking task $\mathcal{T}^{\prime}$. We propose a straightforward exploitation method that is based on the distance in the BEK space. Given an input sample $x_{s}$ and the corresponding BEK vector $\zeta_{h}(x_{s})$ after inference with $h(\cdot)$, the classification of $x_{s}$ for $\mathcal{T}^{\prime}$ is inferred as follows: